Two rectangles have the same width. The length if one is 1 foor longer than the width. The length of the other is 2 feet longer than the width. The larger rectangle has 4 more square feet than the smaller. What is the width of the rectangles?
Answer by ptaylor(2198)


Let x=width
Length of DISABLED_event_one= x+1
Length of other=x+2
Area = Length times Width

Now we are told that:
x(x+2)=x(x+1)+4 clear parens
x^2+2x=x^2+x+4 subtract x^2 and x from both sides
x^2-x^2+2x-x=x^2-x^2+x-x+4 collect like terms
x=4 feet----------------width of the rectangles
x+1=4+1=5 feet--------- length of one
x+2=4+2=6 feet-----------length of the other

ck
Area of smaller=4*5=20 sq ft
Area of larger=4*6=24 sq ft
Larger is 4 more than smaller----checks

Hope this helps ----ptaylor